The science trench becomes very lively again – University of Copenhagen

NEEM > Field diaries folder > UK diaries 2009 > The science trench bec...

23 July 2009

The science trench becomes very lively again

A 1.65 m slab of ice in the line scan instrument in the science trench. The
ice is in one piece with no breaks at all!

A 1.65 m slab of ice in the line scan instrument in the science trench. The ice is in one piece with no breaks at all!

Today was the day where we started processing the ice below the brittle zone. It was a great success. The ice showed no signs of breaking or exploding and we used the day training the team in the science trench. We are ready now – and just need a lot of core. Expectations are high and the whole science team is hoping to be in camp when we pass the transition from interglacial to interglacial (modelled to be at 1390 m). Exchange of people will be next Thursday…

What we have done today:

  1. Drilling with the NEEM long drill: 16.91 m. Drillers depth: 1313.53 m.
  2. Logging 17.36 m of core. Logging depth 1326.82 m.
  3. CFA analysis: 18.15 m. Depth: 500.5 m.
  4. Processing: 8.35 m. Depth: 1289.75 m (bag 2345).
  5. Tuning of saws, alignment of tables and placement of longer guiders in the science trench.
  6. Shallow drilling at the air-firn village. Present depth: 75 m. Air-sampling has been made at the depth 75 m.
  7. Shallow drilling at the second site NEEM2009S1. Depth 88.75 m.
  8. Building the last core buffer in the science trench.

Ad 1: Driller’s Report July 23, 2009:
“Three perfect runs were followed by five very poor runs. Poor runs ended after 60 to 130 cm and were plagued with high current from the start, each ending by a total blockage of drill rotation. Chips transport may be a cause for at least two of these short runs where we found the pump packed and the space between the pump and lowest booster packed, with nothing above the lowest booster. We considered changing the pump, but first we installed new cutters and reduced the pitch and finally a normal run was achieved before dinner without altering the pump or hollow shaft. We will continue to tune the drill during the night shift. Inclination is steady between 2.2 and 2.4 degrees.”

Ad 3. A great celebration just at the time for crew shift at 19:00 when 500 m was passed. CONGRATULAIONS to the CFA team for very dedicated work!

Ad 4. The first bag to be processed was 2331 (start depth: 1281.50 m).

Ad 6: The Laki eruption has been found at the depth 70.98 – 71.08 m by ECM measurements (pls note that the bag system was started wrong with a 1.45 m offset (the first core length!) so depth is: 0.55*BAGNR-1.45). Vas will continue drilling in the S1 hole to a depth of 150 m. The core from 0 to 90 m is Steering Committee core; the core below 88.75 m will be divided as described in the firn-gas program that will be circulated to the SC.

Weather: In the morning sunny but changing to foggy from noon. Night temperatures between Wednesday and Thursday down to -10°C and day temperatures up to -3°C. Variable wind 5-12 kn from SE to SSE. Pressure steadily dropping from 1030 to 1015 mbar.

FL, Dorthe Dahl-Jensen

The football game on the first floor in main dome is very popular and the games become
loads of fun. The ball does not always stay in the box and on the photo
the ball has just been shot out to the left.

The football game on the first floor in main dome is very popular and the games become loads of fun. The ball does not always stay in the box and on the photo the ball has just been shot out to the left.

 

← Previous entry   Next entry →